OpenStudy (bestustar123):

Quadratic equation solving using quadratic formula Write the values of a,b and c for the following equation: x^2+7x-7=0 Then solve using quadratic formula

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (cwrw238):

just compare each term : ax^2 + bx + c = 0 x^2 + 7x - 7 = 0

OpenStudy (cwrw238):

so a = 1 now what is b and c ?

OpenStudy (cwrw238):

note the 1 is understood in x^2 x^2 = 1x^2